What is the 3rd side of triangle if side 1 is 11 ft and side 2 is 9 ft?

Answer by cs1998(1)

The answer is the Square Root of 202... Which is around 14.2126704035519
How we got it. = A^2 +B^2 = C^2
so 9^2 = 81 and 11^2 = 121
81 + 121 = C^2
C^2 = 202
Square Root of 202 = Third side.